(6%) Problem 2: A 0.9 m long section of cable carrying current to a car starter motor makes an angle of 60° with the Earths 5 × 10 T magnetic field Randomized Variables 1=0.9 m B=5×10-5 T F1 = 9.5 × 10-3 N Δ 50% Part (a) What is the current in A when the wire experiences a force of 9.5 x 10-3 N? Grade Summary 0% 100% Potential sin cotan0asin0 atan0acotan s coshOtanhO) cotanh0 tan() acos) sinhO Submissions Attempts remaining: 3 30% per attempt) detailed view 4 5 6 0 END Degrees Radians Submit Hint I give up! Hints: 30% deduction per hint. Hints remaining: 2 Feedback: 30% deduction per feedback -Δ 50% Part (b) If you run the wire through a strong horseshoe magnet, subjecting 5.00 cm of it to a perpendicular 1.75 T field, what force is exerted on the segment of wire in N?
Answer #1

a)

here by using the formla

F = I * L * B sin(theta)

I = 9.5 * 10^-3 / ( 0.9 * 5 * 10^-5 * sin(60)

I = 243.8 amp

the current in the wire 243.8 amp

b)

here we use the same formula

and there sin(theta) = 1 ..... (theta = 90 deg)

F = I * L * B

F = 243.8 * 0.05 * 1.75

F = 21.3 N

the force exerted on the segment of wire is 21.3 N